OpenStudy (anonymous):

find dy/dx assume r, s, and t are constants rx^2 - sy^2 = t^2 i keep on getting ((2t-2xr)/2ys), but it is wrong, any help?

zepdrix (zepdrix):

Ah I see the problem! :)

zepdrix (zepdrix):

Hint: Your final answer shouldn't include any t in it. If t is a constant, then \(\large\rm \frac{d}{dx}t^2=0\) No power rule for him!

OpenStudy (anonymous):

i did (2-2xr)/(2ys) and its still wrong :/

zepdrix (zepdrix):

Yah that doesn't look right. Hmm..

zepdrix (zepdrix):

Did you apply product rule to the (r)(x^2) maybe?

zepdrix (zepdrix):

\[\large\rm r x^2-s y^2=t^2\]Differentiating with respect to x should give you,\[\large\rm 2rx-2syy'=0\]Any confusion on that?

OpenStudy (anonymous):

ok i understand that

zepdrix (zepdrix):

Solve for y' :)

OpenStudy (anonymous):

Yes! that was right, so why was that right though?

zepdrix (zepdrix):

Quote Benitob: `ok i understand that` So clearly you don't understand that derivative if you're asking this question XD lol

zepdrix (zepdrix):

Umm I guess it's because .. chain rule? It's hard to say, I'm not sure what mistakes you were making in your problem.